Acme Aquafers has forecasted the following monthly demand for humidifier sales in the first quarter of 2020:

Month                         Forecasted Demand

January                        14,200 humidifiers

February                      16,800 humidifiers

March                          19,000 humidifiers

Acme has a policy that they have 15% of the following month’s forecasted demand available as finished goods at the end of each month. Given this inventory policy, how many humidifiers must Acme produce in February to support the forecasted sales (assume that January’s ending finished goods inventory were 15% of February’s forecasted demand)?

  1. 2,520 humidifiers
  2. 2,850 humidifiers
  3. 16,800 humidifiers
  4. 17,130 humidifiers
  5. 19,650 humidifiers
Answer #1

Ending Inventory of January month = 16,800 units * 15% = 2,520

Beginning Inventory of February's month will be Ending Inventory of January = 2,520

So, Ending Inventory of February month = 19,000 * 15% = 2,850

Sales in February = 16,800 units

Number of units produced in the month of February = Ending inventory + Sales - Ending Inventory

= 2,850 + 16,800 - 2,520

= 17,130 humidifiers

Option 'D' is correct